assume that the potential energy of a hydrogen atom in its ground state is zero . then its energy in the first excited state will be

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear vikas,

 r_n = {n^2\hbar^2\over Zk_e e^2 m_e}
The smallest possible value of r is called the Bohr radius and is equal to:
r_1 = {\hbar^2 \over (k_e e^2) m_e} = 0.529 \times 10^{-10} \mathrm{m}
The energy of the n-th level is determined by the radius:
 E = -{k_e e^2 \over 2r_n } = - { (k_e e^2)^2 m_e \over 2\hbar^2 n^2} = {-13.6 \mathrm{eV} \over n^2}

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi